OpenStudy (anonymous):

Let f(x) = x2 + 3x – 4 and g(x) = x + 5. Find f(x) • g(x).

OpenStudy (anonymous):

f(x) • g(x) = f(g(x)) f(x+5) = (x+5)^2 + 3(x+5) - 4 = x^2 + 13x + 36

OpenStudy (anonymous):

hope that help you

OpenStudy (anonymous):

I don't think thats right..

OpenStudy (anonymous):

f(x) • g(x) = f(g(x)) f(x+5) = (x+5)^2 + 3(x+5) - 4 = x^2 + 13x + 36 thats what i got